OpenStudy (anonymous):

20x+15x+20(x-1)+15(x-1)=245

OpenStudy (gio333):

x=4

OpenStudy (gio333):

if you want check plug in the 4 where ever you see an x

OpenStudy (anonymous):

add the x's first 20x+15x+20x-20+15x-15=245 70x-35=245 70x=245+35 70x=280 <-- now divide 70 on both sides x=4